9:17イ LTE a campus.college.ch 26 The balance sheet of a business at the start of the week is as folows Freehold premises Funiture and ittings Stock in trade Teade debtors recelvables) 45.000 Capital 3,000 Bank overdaf 28,000 Trade oredtons paya During the week the folowing transactions take piace Stock sold for £11,000cas this stook had cost 000 b Sold stock for C23,000 on credit, this stock had oot £17,000 c) Received cash from trade debtors totalling C18,000 (d) The owners of the business introduced £100,000 ff their own money, which was placed in the business bank acoount e) The owners brought a motor van valued at £10,000 ntthe buses ) Bought stock in trade on credit for £14,000 Paid trade creditors £13,000 Required Show the balance sheet after all of these transactions have been eleced а о
Answer #1
Assets £ Claims £
Freehold premises 145000 Capital 313000
Furniture and fittings 63000 Retained earnings 9000
Motor van 10000 Trade creditors (payables) 24000
Stock in trade 17000
Trade debtors (receivables) 38000
Bank 73000
346000 346000

Working:

Stock in trade = 28000 - 8000 - 17000 + 14000 = $17000

Trade debtors (receivables) = 33000 + 23000 - 18000 = $38000

Bank = -43000 + 11000 + 18000 + 100000 - 13000 = $73000

Capital = 203000 + 100000 + 10000 = $313000

Retained earnings = 11000 - 8000 + 23000 - 17000 = $9000

Trade creditors (payables) = 23000 + 14000 - 13000 = $24000